OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = -33 and a9 = -121? What is the sum of a 22-term arithmetic sequence where the first term is 57 and the last term is -27?

OpenStudy (campbell_st):

1st task is to find the common difference use theformula for a term in an arthmetic sequence\[T _{n}=a+(n-1)d\] use term 9 \[-121 = -33 +(9 -1)d\] -121 = -33 + 8d -88 = 8d d = 8 the common difference for the sequence is 8 use the above formula to find the 32 term\[T _{32} = -33 + (32-1)\times(-8)\] you can evaluate it sum of an arthmetic series is \[s _{n}=\frac{n}{2}(a + l)\] n = number of terms a = 1st term l = last term

OpenStudy (anonymous):

I got -281 but that is not one of my options.

OpenStudy (campbell_st):

for the sum of the term

OpenStudy (campbell_st):

oops calculation error the common difference is d = -11 and not 8

OpenStudy (campbell_st):

-88 = 8d then d = -11